# Lesser Green’s Function and Chirality Entanglement Entropy via the In-Medium NJL Model

### v1

@@ Abstract
We study chiral symmetry restoration in hot-and-dense quark matter using the von Neumann chirality entropy within the in-medium Nambu-Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$, the reduced correlator $C_{L}=P_{L}G^{<}P_{L}$ is constructed, and the associated entropy $S_{\chi}=-\mathrm{Tr}[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})]$ is introduced to quantify the entanglement between left- and right-handed quark sectors. The dynamical quark mass $M_{q}(T,\mu_{q})$ obtained from the gap equation exhibits the expected QCD-like phase structure: A second-order transition in the chiral limit and a smooth crossover for finite $m_{q}$. The chirality entropy $S_{\chi}$ increases monotonically with temperature and chemical potential, approaching a maximal value as $M_{q}\to 0$. We also explore the critical exponents and scaling behavior of those quantities, yielding $\beta_{S_{\chi}}\simeq 1$. This demonstrates that $S_{\chi}$ serves as an information-theoretic probe for chiral symmetry restoration, linking dynamical mass generation to quantum entanglement in strongly interacting matter.
@@ I Introduction
The spontaneous breakdown (SBCS) and restoration of chiral symmetry constitute a fundamental aspect of Quantum Chromodynamics (QCD) and play a central role in the mass generation mechanism of hadrons and in the formation of the quark-gluon plasma (QGP) at high temperature or density [1, 2, 3, 4, 5]. In the QCD vacuum, chiral symmetry is dynamically broken, leading to a large constituent quark mass and giving rise to light pseudoscalar mesons as Nambu-Goldstone bosons. As temperature or baryon density increases, the quark condensate $\langle\bar{q}q\rangle$ melts, the constituent mass decreases, and chiral symmetry is partially restored. Lattice QCD simulations have established that this restoration occurs through a smooth crossover at vanishing chemical potential for physical quark masses [6, 7]. At the same time, effective models such as the Nambu–Jona-Lasinio (NJL), Polyakov-NJL (PNJL), and instanton-based models successfully reproduce the qualitative features of the QCD phase diagram [2, 3, 8, 9, 10].
Recently, concepts from quantum information theory have been applied to strongly interacting systems to characterize their microscopic coherence and correlations in terms of entanglement entropy [11, 12, 13, 14]. In QCD and other field-theoretical systems, entanglement-based observables have been used to probe confinement, deconfinement, topological ordering, and spin structures [15, 16, 17, 18]. The von Neumann entropy, in particular, provides a quantitative measure of quantum coherence loss and has been proposed as a diagnostic tool for phase transitions in gauge theories and many-body systems.
In this work, we formulate the von Neumann chirality entropy as an information-theoretic probe of chiral symmetry restoration within the in-medium NJL model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism [19, 20], we construct the reduced left-handed correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and define the entropy
@@
This entropy measures the degree of quantum entanglement between left- and right-handed quark sectors, which is dynamically generated through the mass term $M_{q}(\bar{q}_{L}q_{R}+\bar{q}_{R}q_{L})$. In the chirally broken phase ($M_{q}\neq 0$), the left and right components are polarized and coherence is maintained, while in the restored phase ($M_{q}\to 0$), they become maximally mixed, leading to a large $S_{\chi}$.
We demonstrate that $S_{\chi}(T,\mu_{q})$ increases monotonically with both temperature and chemical potential, reflecting the progressive loss of chiral coherence and the transition to a maximally mixed state. The resulting behavior mirrors the thermodynamic phase structure of the NJL model: A second-order transition in the chiral limit and a smooth crossover for finite current quark mass. Thus, the von Neumann chirality entropy serves as an information-theoretic probe that connects the dynamical mechanism of mass generation to the microscopic structure of quantum entanglement in QCD matter.
Most existing entanglement-related studies in QCD address spatial bipartition entropy, Rényi entropies, or color entanglement in Euclidean or lattice formulations. In contrast, the present work develops an information-theoretic quantity associated with the chiral internal degree of freedom, constructed directly from the real-time lesser Green function. The resulting chirality-reduced entropy captures dynamical changes in the quasiparticle spectrum and offers a new, complementary perspective on the chiral phase transition beyond conventional order parameters and previously studied entanglement measures.
The remainder of this paper is organized as follows. In Sec. II, we outline the theoretical framework of the finite-temperature Nambu-Jona-Lasinio model and introduce the lesser Green’s function in the real-time formalism, which provides the foundation for the subsequent definition of the reduced correlator. In Sec. III, we derive the left-handed reduced correlator $C_{L}(\bm{k})$ and analyze its connection to the dynamical mass and chiral mixing. In Sec. IV, the von Neumann chirality entropy is formulated and its physical interpretation as an information-theoretic measure of chiral entanglement is discussed in detail. Section V presents the numerical results for the dynamical quark mass and chirality entropy as functions of temperature and chemical potential, together with a detailed physical discussion of the QCD-like phase structure. Finally, Sec. VI summarizes the main conclusions and remarks on possible extensions of the present approach, including applications to the PNJL model and to lattice-QCD analyses of the entanglement in hadronic matter.
@@ VII Summary and Outlook
In this work, we have developed a consistent framework to quantify the quantum entanglement between left- and right-handed quark sectors within the finite-temperature and density Nambu–Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism, we derived the left-handed reduced correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and introduced the von Neumann chirality entropy, $S_{\chi}=-\mathrm{Tr}\!\left[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})\right]$, as an information-theoretic measure of chiral coherence, this entropy characterizes the degree of mixing between the two chiral components, thereby providing a direct link between dynamical mass generation and the quantum information content of QCD matter.
@@
The self-consistent solutions of the NJL gap equation reveal a QCD-like chiral phase structure: A second-order transition in the chiral limit and a smooth crossover for finite current quark mass. The dynamical quark mass $M_{q}(T,\mu_{q})$ decreases continuously with temperature and chemical potential, signaling progressive chiral symmetry restoration. Using this behavior, we computed the chirality entanglement entropy density $S_{\chi}/V$ and found that it increases monotonically with both $T$ and $\mu_{q}$, approaching a maximal value as $M_{q}\to 0$. Thus, $S_{\chi}$ mirrors the order-parameter behavior of $M_{q}$, i.e., it is nearly zero in the broken phase, where the system is chirally polarized, and maximal in the restored phase, where left- and right-handed components are thoroughly mixed.
@@
A key aspect of this work is the scaling analysis of $S_{\chi}$ near the second-order critical line in the chiral limit. The dynamical mass exhibits the mean-field behavior $\beta_{M_{q}}\simeq 1/2$, being consistent with the standard Landau universality of the NJL model. In contrast, the entropy provides $\beta_{S_{\chi}}\simeq 1$, i.e., a distinctly larger exponent. This difference arises because $S_{\chi}$ is the temperature derivative of the free energy. At the same time, the critical part of the latter is dominated by the order-parameter amplitude squared, $\Omega_{\rm crit}\sim M_{q}^{2}\sim(T_{c}-T)^{1}$. Consequently, the entropy inherits the $M_{q}^{2}$ scaling rather than $M_{q}$, leading to an exponent approximately twice that of the order parameter. This confirms that $S_{\chi}$ is not an order parameter itself, but a thermodynamic entanglement response that nevertheless remains tightly correlated with the underlying critical dynamics.
@@
The monotonic increase of $S_{\chi}(T,\mu_{q})$ therefore encodes the loss of left-right quantum coherence as the system evolves from a strongly correlated vacuum to a thermally randomized quark ensemble. In this sense, $S_{\chi}$ serves as an information-theoretic probe for chiral symmetry restoration, complementing the traditional condensate $\langle\bar{q}q\rangle$ by emphasizing the quantum entanglement structure of the quark sector. The framework developed here thus bridges thermal quantum field theory and quantum information theory, allowing symmetry restoration to be interpreted as the emergence of maximal quantum mixing in the chiral subspace.
@@
The present study opens several directions for future work. Incorporating the Polyakov loop or explicit color interactions would enable a combined analysis of deconfinement and chirality entanglement. Extensions to PNJL or Dyson-Schwinger frameworks could clarify the interplay among confinement, dynamical mass, and information entropy. In confining models such as the PNJL framework, color-singlet constraints suppress single-quark excitations below the deconfinement temperature, reducing the accessible Hilbert-space volume and lowering $S_{\chi}$ in the confined phase. This suggests that confinement delays the onset of large chirality entanglement until color screening becomes effective near $T_{c}$. Future lattice-QCD simulations may directly evaluate fermionic correlation matrices or reduced chiral projectors to obtain $S_{\chi}$ nonperturbatively and thus test our entanglement-based interpretation of chiral symmetry restoration, since $C_{L}(\bm{k})$ is constructed from fermionic two-point functions and the same reduced correlator can in principle be evaluated on the lattice. Such developments would further establish the von Neumann chirality entropy as a universal, information-based diagnostic of phase transitions in strongly interacting quantum systems. Related works will appear elsewhere.

### v2

@@ Abstract
We study chiral symmetry restoration in hot and dense quark matter using the von Neumann chirality entropy within the in-medium Nambu–Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$, we construct the chirality-reduced correlator $C_{L}=P_{L}G^{<}P_{L}$ and define the associated entropy $S_{\chi}=-\mathrm{Tr}\left[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})\right]$ to quantify quantum entanglement between left- and right-handed quark sectors. The dynamical quark mass $M_{q}(T,\mu_{q})$ reproduces the expected QCD-like phase structure, showing a second-order transition in the chiral limit and a smooth crossover for finite current quark mass. The chirality entropy $S_{\chi}$ increases monotonically with temperature and chemical potential and approaches a maximal value as $M_{q}\to 0$. Analyzing its critical behavior, we find a scaling exponent $\beta_{S_{\chi}}\simeq 1$, distinct from that of the chiral order parameter. This indicates that $S_{\chi}$ is not an order parameter but a thermodynamic measure of chiral quantum decoherence. Our results demonstrate that chiral symmetry restoration and chiral decoherence are not identical phenomena, and that the chirality entropy reveals information inaccessible to conventional symmetry-breaking observables.
@@ I Introduction
The spontaneous breakdown (SBCS) and restoration of chiral symmetry constitute a fundamental aspect of Quantum Chromodynamics (QCD) and play a central role in the mass generation mechanism of hadrons and in the formation of the quark-gluon plasma (QGP) at high temperature or density [1, 2, 3, 4, 5]. In the QCD vacuum, chiral symmetry is dynamically broken, leading to a large constituent quark mass and the emergence of light pseudoscalar mesons as Nambu-Goldstone bosons. As temperature or baryon density increases, the quark condensate $\langle\bar{q}q\rangle$ melts, the constituent mass decreases, and chiral symmetry is partially restored. Lattice QCD simulations have established that this restoration occurs through a smooth crossover at vanishing chemical potential for physical quark masses [6, 7]. At the same time, effective models such as the Nambu–Jona-Lasinio (NJL), Polyakov-NJL (PNJL), and instanton-based models successfully reproduce the qualitative features of the QCD phase diagram [2, 3, 8, 9, 10].
Recently, concepts from quantum information theory have been applied to strongly interacting systems to characterize microscopic coherence and correlations in terms of entanglement entropy [11, 12, 13, 14]. In QCD and related field-theoretical systems, entanglement-based observables have been explored as probes of confinement, deconfinement, topological ordering, and spin structures [15, 16, 17, 18]. In particular, the von Neumann entropy provides a quantitative measure of quantum mixedness and coherence loss, offering a complementary diagnostic to conventional symmetry-based observables.
In this work, we formulate the von Neumann chirality entropy as a probe of chiral symmetry restoration within the in-medium NJL model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism [19, 20], we construct the chirality-reduced correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and define the entropy
@@
This quantity measures the quantum entanglement between left- and right-handed quark sectors generated by the mass term $M_{q}(\bar{q}_{L}q_{R}+\bar{q}_{R}q_{L})$. In the chirally broken phase ($M_{q}\neq 0$), the two chiral components remain polarized and coherence is largely preserved, whereas in the restored phase ($M_{q}\to 0$) they become maximally mixed, leading to an enhanced $S_{\chi}$.
We demonstrate that $S_{\chi}(T,\mu_{q})$ increases monotonically with both temperature and chemical potential, reflecting the progressive loss of chiral coherence. This behavior follows the thermodynamic phase structure of the NJL model, exhibiting a second-order transition in the chiral limit and a smooth crossover for finite current quark mass. While this trend parallels that of the dynamical mass, $S_{\chi}$ characterizes the evolution of quantum mixing rather than symmetry breaking itself.
It is important to emphasize that the chirality entropy introduced here is conceptually distinct from the conventional chiral order parameter $\langle\bar{q}q\rangle$. The quark condensate is a one-point expectation value that quantifies the magnitude of symmetry breaking, whereas $S_{\chi}$ is defined through the spectrum of a reduced fermionic correlation matrix obtained by tracing out the opposite chiral sector. As such, $S_{\chi}$ directly measures the loss of purity of a chiral subsystem and provides an information-theoretic characterization of chiral mixing. Although correlated with the dynamical mass in mean-field descriptions, $S_{\chi}$ is not reducible to $\langle\bar{q}q\rangle$, as evidenced by its distinct critical behavior, with scaling exponents $\beta_{S_{\chi}}\simeq 1$ and $\beta_{M_{q}}\simeq 1/2$ near the chiral transition.
Most existing entanglement-related studies in QCD focus on spatial bipartite entropy, Rényi entropies, or color entanglement in Euclidean or lattice formulations. In contrast, the present work introduces an entanglement measure associated with an internal chiral degree of freedom, constructed directly from the real-time lesser Green’s function. The resulting chirality-resolved entropy captures dynamical changes in the quasiparticle spectrum and provides a complementary perspective on the chiral phase transition beyond conventional order parameters.
The remainder of this paper is organized as follows. In Sec. II, we outline the finite-temperature NJL framework and the real-time formalism for the lesser Green’s function. In Sec. III, we derive the reduced correlator $C_{L}(\bm{k})$ and analyze its relation to dynamical mass generation and chiral mixing. Section IV introduces the von Neumann chirality entropy and discusses its physical interpretation. Numerical results for the dynamical mass and chirality entropy are presented in Sec. V. Finally, Sec. VI summarizes the main conclusions and discusses possible extensions, including applications to PNJL models and lattice-QCD studies of chiral entanglement.
@@ V Von Neumann Entropy and Chirality Entanglement
We stress that the chirality entropy $S_{\chi}$ cannot be expressed as a functional of the chiral condensate $\langle\bar{q}q\rangle$ alone. The condensate probes the expectation value of the scalar bilinear, whereas $S_{\chi}$ depends on the full spectrum of the reduced correlator $C_{L}$ and therefore on the momentum-resolved occupation probabilities $\nu_{k}$. As a result, $S_{\chi}$ captures information about quantum mixedness that is invisible to local order parameters.
@@ VI Numerical Results and Discussions
An important advantage of the entanglement-based description is that the entropy derivative $dS_{\chi}/dT$ acts as an entanglement susceptibility. As shown in Fig. 4, the peak of $dS_{\chi}/dT$ appears at a lower temperature than that of $|dM_{q}/dT|$, indicating that chiral quantum decoherence sets in before the rapid melting of the dynamical mass. This separation cannot be inferred from the condensate alone and highlights the additional physical information contained in $S_{\chi}$. This clearly demonstrates that the chirality entropy reveals information inaccessible to symmetry-breaking observables, capturing the onset of chiral decoherence beyond conventional order-parameter diagnostics.
@@
The distinct critical exponent $\beta_{S_{\chi}}\simeq 1$ confirms that the chirality entropy is not an order parameter. Instead, it reflects the thermodynamic response of quantum entanglement to the underlying critical dynamics. This derivative-type scaling behavior is absent in the chiral condensate and demonstrates that $S_{\chi}$ encodes complementary information beyond conventional symmetry-breaking observables.
@@ VII Summary and Outlook
Our results demonstrate that chiral symmetry restoration and chiral decoherence are not identical phenomena, even in mean-field QCD-like models. In this work, we have developed a consistent framework to quantify quantum entanglement between left- and right-handed quark sectors within the finite-temperature and density Nambu–Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism, we derived the left-handed reduced correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and introduced the von Neumann chirality entropy $S_{\chi}=-\mathrm{Tr}\!\left[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})\right]$ as an information-theoretic measure of chiral coherence. This entropy quantifies the degree of quantum mixing between the two chiral components and establishes a direct connection between dynamical mass generation and the information content of QCD matter. We emphasize that the chirality entanglement entropy is not a reformulation of the chiral condensate, but a distinct diagnostic of quantum coherence loss between chiral sectors.
@@
The self-consistent solutions of the NJL gap equation reproduce a QCD-like chiral phase structure, featuring a second-order transition in the chiral limit and a smooth crossover for finite current quark mass. The dynamical quark mass $M_{q}(T,\mu_{q})$ decreases continuously with increasing temperature and chemical potential, signaling progressive chiral symmetry restoration. Using these solutions, we computed the chirality entanglement entropy density $S_{\chi}/V$ and found that it increases monotonically with both $T$ and $\mu_{q}$, approaching a maximal value as $M_{q}\to 0$. While $S_{\chi}$ follows the overall thermodynamic trend of the order parameter, it characterizes the evolution of quantum mixing rather than symmetry breaking itself.
@@
A central result of this work is the scaling behavior of $S_{\chi}$ near the second-order critical line in the chiral limit. The dynamical mass exhibits the expected mean-field exponent $\beta_{M_{q}}\simeq 1/2$, consistent with Landau universality. In contrast, the chirality entropy shows a distinct exponent $\beta_{S_{\chi}}\simeq 1$. This difference originates from the thermodynamic structure of the free energy: since $S_{\chi}$ corresponds to a temperature derivative and the critical contribution satisfies $\Omega_{\rm crit}\sim M_{q}^{2}\sim(T_{c}-T)^{1}$, the entropy inherits the $M_{q}^{2}$ scaling rather than that of the order parameter itself. This confirms that $S_{\chi}$ is not an order parameter, but a thermodynamic response encoding the growth of chiral quantum mixing.
@@
The monotonic increase of $S_{\chi}(T,\mu_{q})$ therefore reflects the progressive loss of left-right quantum coherence as the system evolves from a chirally polarized vacuum to a thermally randomized quark ensemble. In this sense, $S_{\chi}$ serves as an information-theoretic probe of chiral symmetry restoration, complementing conventional symmetry-breaking observables by accessing the entanglement structure of the quark sector. Chiral restoration may thus be interpreted not only as the melting of an order parameter, but as the emergence of maximal quantum mixing in the chiral subspace.
@@
The present study opens several directions for future work. Incorporating the Polyakov loop or explicit color interactions would enable a combined analysis of deconfinement and chirality entanglement. In confining extensions such as the PNJL framework, color-singlet constraints suppress single-quark excitations below the deconfinement temperature, reducing the accessible Hilbert-space volume and lowering $S_{\chi}$ in the confined phase. This suggests that confinement delays the onset of large chirality entanglement until color screening becomes effective near $T_{c}$. The temperature offset $\Delta T\simeq 18$ MeV between the peaks of $dS_{\chi}/dT$ and $|dM_{q}/dT|$ observed in Fig. 4 provides an experimentally testable consequence: observables sensitive to quantum coherence, such as fluctuation-based measures or two-particle interferometry in heavy-ion collisions, may signal the onset of chiral restoration at lower temperatures than bulk thermodynamic quantities.
Future lattice-QCD simulations may directly evaluate fermionic correlation matrices or chirality-projected correlators to obtain $S_{\chi}$ nonperturbatively, since $C_{L}(\bm{k})$ is constructed from fermionic two-point functions accessible on the lattice. Such studies would provide a direct test of the entanglement-based interpretation of chiral symmetry restoration proposed here and further establish the von Neumann chirality entropy as an information-based diagnostic of phase transitions in strongly interacting matter.